2. Please test hypothesis for a mean or for a proportion. Please state null and alternative hypothesis

Test Hypotheses of population percent = 33%, sample percent = 31%,n= 120, confidence level = 99%
Answer #1

H0: p = 0.33

Ha: p 0.33

Test statistics

z = ( - p) / sqrt [ p( 1 - p) / n ]

= ( 0.31 - 0.33) / sqrt [ 0.33 ( 1 - 0.33) / 120 ]

= -0.47

p-value = 2 * P(Z < z)

= 2 * P(Z < -0.47)

= 2 * 0.3192

= 0.6384

Since p-value > 0.01 level, Fail to reject H0.